Given the system of equations, what is the value of the determinant?
3x + y = 9
2x - 4y = 6

Answer by Alan3354(69443)

Given the system of equations, what is the value of the determinant?
3x + y = 9
2x - 4y = 6
------------
-12 - 2 = -14
Det = -14